OpenStudy (anonymous):

The expression 4n – 1 is a rule to describe a certain sequence. What are the first three terms of the sequence?

OpenStudy (anonymous):

starting from 0?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

4(0)-1=-1 4(1)-1=3 4(2)-1=7

OpenStudy (anonymous):

thanks alot

OpenStudy (anonymous):

thank you!